OpenStudy (anonymous):

A school is ordering boxes of copy paper for its teachers. When placing their order, the paper company explains the pricing to them as follows. If they order no more than 1000 boxes, the company will charge them $35 per box. If they order more than 1000 boxes, the company will charge them $27.50 per box. Which type of function would best model this situation: a step function or a piecewise function that is not a step function? Explain.

OpenStudy (anonymous):

step function

OpenStudy (anonymous):

27.5would be better

OpenStudy (anonymous):

Remember me?

OpenStudy (anonymous):

@AnswerMyQuestions yes i do we're all back again

OpenStudy (anonymous):

@AnswerMyQuestions oh my gosh yessss

OpenStudy (anonymous):

@HARSH123 can you explain it plz

OpenStudy (anonymous):

3 people on this post that were on abby's looooooooooooooooooooong post.

OpenStudy (anonymous):

yes it was soooooo long

OpenStudy (anonymous):

@AnswerMyQuestions very long

OpenStudy (anonymous):



OpenStudy (anonymous):

@AnswerMyQuestions that was about a stupid guy she thought wanted her

OpenStudy (anonymous):

ok This is a step function since it has a finite number of piece

OpenStudy (anonymous):

f(x) = 35*x if x < or = 1000 = 27.50*x if x > 1000

OpenStudy (anonymous):

@HARSH123 thanks

OpenStudy (anonymous):

this function has two pieces and therefore is a step function \and we can can that it is step function is a piecewise constant function having only finitely many pieces.

OpenStudy (anonymous):

ok my pleasure

OpenStudy (anonymous):

HEYYYY

OpenStudy (anonymous):

...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@AnswerMyQuestions yup swag